OpenStudy (anonymous):

Multiply. (x+3)/(2x-1) times (x-1)/(x+3)

OpenStudy (anonymous):

\[\frac{ (x+3) }{ (2x-1) } \times \frac{ (x-1) }{ (x+3) }\]

OpenStudy (fifciol):

e.g. if you multiply\[\frac{ 6 }{ 5}*\frac{ 1 }{ 6} \] what can you do with those six?